### v1

# Exceptional topology in Non-Hermitian Twisted Bilayer Graphene

Abstract
Twisted bilayer graphene has extraordinary electronic properties at the magic angle along with an isolated flat band at magic angle. However, the non-Hermitian phenomena in twisted bilayer graphene remain unexplored. In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss. Using a non-Hermitian generalization of Bistritzer-MacDonald(BM) model, we find Dirac cones centered at corner of the moiré Brillouin zone deformed in the presence of non-Hermiticity. This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both and corners of the Brillouion zone. The coincident of exceptional rings at point characterizes an “exceptional magic angle”, at which the system hosts flat bands with zero energy and finite lifetime. More interestingly, we find that the topological charge in the moiré Brillouin zone is conserved during the expansion and fusion of the exceptional ring, which is absent in two-dimensional systems constraining by Nielsen-Ninomiya theorem. These findings can be demonstrated in realistic cold atom and metamaterial systems and will stimulate further study on non-Hermitian phenomena in twistronic.
I introduction
Graphene owes extraordinary electronic and optical properties in two-dimensional (2D) systems [1, 2]. These properties relate to its lattice structure, which is directly interconnected with its wave-function topology. In single-layer graphene, the presence of unusual linear band dispersion termed Dirac band dispersion at and points of the Brillouin zone is derived from its honeycomb lattice geometry. The Lorentz invariant Dirac cone at and points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4]. When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6]. One prominent example is a graphene bilayer with a relatively small angle rotation between the layers [7]. The recent studies on twisted bilayer graphene (TBG) focus on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12]. It is related to a variety of exotic phrases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29].
The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [30, 31, 32]. The presence of non-Hermiticity can transform Hermitian degenerate points, such as a Dirac-like point or Weyl point, into a ring of exceptional points [33, 34]. At an exceptional point, not only the real and imaginary parts of the eigenvalues degenerate, but also the eigenvectors coalesce. Many phenomena are unique to non-Hermitian systems, including non-Hermitian skin effect [35] and non-Hermitian topological classifications [36, 37].
Despite topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of twisted moiré system with magic angle has been overlooked [38, 39].
In particular, twisted bilayer graphene has extraordinary topological properties. Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in 2D periodic systems following the Nielsen-Ninomiya theorem [40, 41]. This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges. Actually, when the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [34]. It is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
In this paper, we uncover a new type of twisted band engineering in TBG, with non-Hermitian perturbation on one of the layers. As shown in Fig. 1(a), balanced gain (in bright color) and loss (in light color) are put on the bottom layers (in red and blue, respectively). The superposing of two graphene layers with a twist angle between them creates moiré patterns. In the mBZ shown in Fig. 1(b), the low-energy band structure consists of Dirac cones from the rotated layers located at the and corners. When the non-Hermitian perturbation is turned on the bottom layer, Dirac points morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig. 1(c). We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at points in an mBZ. We call this phenomenon “exceptional magic”. And importantly, our results demonstrated that the total Berry charge is conserved in the mBZ even after the merging of exceptional rings, this is a large difference from non-Hermitian single-layer graphene.
The paper is organized as follows. Section II introduces the non-Hermitian TBG model. Both the real and imaginary parts of the low-energy moiré band spectrum will be explored in Sec. III. In Sec. IV, we will use a non-Hermitian three triod model to discuss exceptional physics. In Sec. V, we calculate Beery curvature and the corresponding Chern number. Finally, we conclude our results in Sec. VI.
II The non-Hermitian bilayer model
We consider balanced gain and loss on both layers of the bilayer system. We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene. The low-energy Hamiltonian for the mBZ is:
where the top layer Hamiltonian is
and the bottom layer Hamiltonian is non-Hermitian
Here, is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity. And corresponds to the rotated Pauli matrices . is the sublattice-dependent moiré potential that couples the two layers. The interlayer coupling is in the form
with and the interlayer coupling strength in the AA (AB) region. The unit vectors , . is the moiré modulation vector and is the magnitude of the Dirac wave vector, where is the lattice constant of monolayer graphene. The hopping strength encodes the inter-layer coupling in the AB region.
The generators of the magnetic space group for the Hermitian twisted bilayer graphene include a rotation symmetry, a rotation symmetry, and a symmetry [42]. The single-valley Hamiltonian only has the symmetry, since and symmetries play the role of mapping one valley to the other. We can check that the non-Hermitian model preserves the rotation symmetry.
If there is no AA stacking (), the Hamiltonian preserves chiral symmetry under chiral symmetry operator in the absence of non-Hermiticity [46, 43, 44, 42].
In the presence of the balanced gain and loss term, the chiral symmetry is broken but non-Hermitian chiral symmetry is preserved In the following numerical calculation, we take , = 110meV and . They correspond to the first magic angle .
III The appearance and development of exceptional rings
Let us investigate the band structure for twist angle . In the absence of non-Hermiticity, this is not a magic angle with a flat band. If the interlayer coupling effect is absent (), the dispersion in the moiré Brillouin zone is equivalent to the folded dispersion in the Brillouin zone of graphene. The Dirac point of the point in layer 1 is mapped to , while that of the point of layer 2 is mapped to .
In the presence of non-Hermitian perturbation , the energy dispersion becomes complex. In Fig. 2(a), we can see that non-Hermitian perturbation deforms the Dirac cones at corners. More specifically, the degeneracy point at the points in the real part of the spectrum morphs into a ring of degeneracy points. This is a stark difference from single-layer graphene with onsite gain and loss. The latter has deformed Dirac cones in both and corners in the Brillouin zone. This difference is due to the fact that and in non-Hermitian TBG are folded from different graphene layers. The non-Hermitian can only deform the Dirac cone on the bottom layer.
In particular, the degeneracy of the two lowest energy states leads to the appearance of singularity points in momentum space. At the same time, a ring of degeneracy point appears at the same momentum position in the imaginary part of the spectrum, as shown in Fig. 2(b). We found that not only the eigenvalues coalesce but also their corresponding eigenvectors coalesce at this point. This confirms that these singularity points are exceptional points in momentum space. Thus, the Dirac cone is turned into a ring of exceptional points, this is similar to that in non-Hermitian semimetals [33].
At a critical non-Hermitian strength, the exceptional point approach point, where the whole lowest band becomes flat, as shown in Fig. 3. Different metric including the vanishing of Dirac velocity, maximum gap to neighboring bands, and minimum bandwidth gives the same results. The value of this critical non-Hermiticity strength depends on the twist angle .
IV Exceptional cone
To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. LABEL:eq:H at the first honeycomb shell. The Hamiltonian is written as
where for are on the top layer and on the bottom layer. Since is non-Hermitian, the Hamiltonian has biorthogonal eigenvectors. In this system, the left and right eigenvectors are four two-component spinors
Using perturbation theory one can derived the effective Hamiltonian fin the space of The Hamiltonian can be divide into a momentum-independent part and a momentum-dependent part . The can be expressed by
Since is independent of non-Hermicity, we can expect , which is similar to the Hermitian model. Thus, the spinor in the zero-energy eigenstates satisfies
This form seems to be the same as that of the Hermitian TBG. However, is in a non-Hermitian form, the is different from . We can check that the biorthorgonal wave functions are normalized as
The effective Hamiltonian matrix to leading order in is
with the renormalized velocity being
and the renormalized non-Hermitian strength being
We can see that aside from a renormalized velocity and non-Hermitian strength, the form of Hamiltonian is identical to the continuum model Hamiltonian of non-Hermitian single-layer graphene. The form of the renormalized velocity is the same as that of the Hermitian TBG. It decreases with increasing interlayer coupling in AB region . To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases. When , increases with increasing interlayer coupling in AB region . Thus, the ratio becomes larger as increases. When, increases with increasing interlayer coupling in AA region . The enon-Hermitian effect is more complicated.
V Berry Curvature and Chern number
So far, we have shown the existence of exceptional geometries in the non-Hermitian TBG. In this section, we will explore their topological properties numerically.
The topological charge or Chern number of an exceptional contour can be obtained by integrating the Berry connection along a closed surface containing the exceptional contour
where the Berry connection is defined by the eigenvectors as
Here, and can be the left or right eigenvectors. Since previous study has proven that the total topological charge is the same for four possible choices of eigenvectors [47], we will choose and in the following calculation.
In numerical calculation, the Berry connection is presented by the local Berry curvature , which can be calculated as
with being half of the lattice constant.
Figure 4 shows the Berry curvature in mBZ. We can see that Berry curvatures have sharp peaks around high symmetric lines, which are close to the boundary of exceptional rings. This can be well understood since Berry curvature is the phase of an overlap tracing a small loop in mBZ. On the boundary of exceptional rings, the wave functions vary rapidly and will give a large Berry curvature.
Upon summing the Berry curvature on the mmBZ, we can obtain the total Berry charge or Chern number
The non-Hermitian perturbation transforms a Dirac point into an exceptional ring. In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners. We find that the total Berry charge in mBZ is 1 throughout the merging of exceptional rings with increasing (e.g., for ). At much larger , the Berry charge can be a quantized value larger than 1 due to approaching band crossings with the remote bands. This supports our expectation that the topological charge in the system is preserved in the exceptional contour. And most importantly, this characterizes the topological features of non-Hermitian TBG. This is opposite to the merging of exceptional contours with opposite topological charge, in which the topological charge is dissipated and gives a single, uncharged exceptional contour.
VI Discussion and conclusion
In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers. We found that exceptional rings centered at corner in the moiré Brillouin zone in the () valley. This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene. The exceptional rings coincide and form bands whose real parts are flat at a specific non-Hermitian strength, which is denoted as “exceptional Dirac magic”. This flat band is characterized with a robust Chern number in moiré Brillouin zone. This is different from single-layer graphene in which Dirac/Weyl exceptional rings carrying opposite topological charges are coupled and annihilated.
We have introduced a new design concept for twisted moiré systems, in which a certain “exceptional Dirac magic” angle is achieved. Non-Hermitian TBG can be realized in cold atom and metamaterial system. On one hand, due to the great advantage of manipulating atom-atom interaction and light-matter coupling, ultracold atom experiments provide an unprecedented opportunity. Remarkably, atomic Bose-Einstein condensate has been recently realized in twisted bilayer optical lattices [57]. Also, there are proposals on the photonic analog [48, 49, 50, 51, 52] and phononic analog [53, 54] of twisted bilayer graphene. In particular, there is demonstration of optical bilayer photonic crystal devices in the microwave range [55] and in the optical frequency range [56]On the other hand, on-site dissipations are well-controlled non-Hermitian terms that can be realized experimentally in many systems [30, 31, 32]. In particular, the Weyl exceptional ring has been realized in photonic experiment [58, 59].
The highlight of this work is the discovery of non-dissipation exceptional topology. The interplay of topology, non-Hermitian, and moiré physics can open the door to new phenomena. Especially, the reduced kinetic energy of the flat bands in twisted bilayer graphene makes it a correlated system. Non-Hermitian strongly interacting Dirac system has recently been calculated [60]. The non-Hermitian physics in quantum many-body systems, such as strongly correlated phases in moié system, will be a future focus.
Note added-. Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 12104099 and No. 12274095) and Guangzhou Science and Technology Program (Grant No. 2024A04J0272)References
- Neto et al. [2009] A. C. Neto, F. Guinea, N. M. Peres, K. S. Novoselov, and A. K. Geim, The electronic properties of graphene, Reviews of modern physics 81, 109 (2009).
- Gonçalves and Peres [2016] P. A. D. Gonçalves and N. M. Peres, An introduction to graphene plasmonics (World Scientific, 2016).
- Novoselov et al. [2005] K. S. Novoselov, A. K. Geim, S. V. Morozov, D. Jiang, M. I. Katsnelson, I. V. Grigorieva, S. V. Dubonos, and A. A. Firsov, Two-dimensional gas of massless dirac fermions in graphene, nature 438, 197 (2005).
- Zhang et al. [2005] Y. Zhang, Y.-W. Tan, H. L. Stormer, and P. Kim, Experimental observation of the quantum hall effect and berry’s phase in graphene, nature 438, 201 (2005).
- Charlier et al. [1992] J.-C. Charlier, J.-P. Michenaud, and P. Lambin, Tight-binding density of electronic states of pregraphitic carbon, Phys. Rev. B 46, 4540 (1992).
- McCann [2006] E. McCann, Asymmetry gap in the electronic band structure of bilayer graphene, Phys. Rev. B 74, 161403 (2006).
- Dos Santos et al. [2007] J. L. Dos Santos, N. Peres, and A. C. Neto, Graphene bilayer with a twist: electronic structure, Physical review letters 99, 256802 (2007).
- Bistritzer and MacDonald [2011a] R. Bistritzer and A. H. MacDonald, Moiré bands in twisted double-layer graphene, Proceedings of the National Academy of Sciences 108, 12233 (2011a).
- Bistritzer and MacDonald [2011b] R. Bistritzer and A. MacDonald, Moiré butterflies in twisted bilayer graphene, Physical Review B 84, 035440 (2011b).
- Dos Santos et al. [2012] J. L. Dos Santos, N. Peres, and A. C. Neto, Continuum model of the twisted graphene bilayer, Physical Review B 86, 155449 (2012).
- Shallcross et al. [2010] S. Shallcross, S. Sharma, E. Kandelaki, and O. A. Pankratov, Electronic structure of turbostratic graphene, Phys. Rev. B 81, 165105 (2010).
- Luican et al. [2011] A. Luican, G. Li, A. Reina, J. Kong, R. Nair, K. S. Novoselov, A. K. Geim, and E. Andrei, Single-layer behavior and its breakdown in twisted graphene layers, Physical review letters 106, 126802 (2011).
- Serlin et al. [2020] M. Serlin, C. Tschirhart, H. Polshyn, Y. Zhang, J. Zhu, K. Watanabe, T. Taniguchi, L. Balents, and A. Young, Intrinsic quantized anomalous hall effect in a moiré heterostructure, Science 367, 900 (2020).
- Sharpe et al. [2019] A. L. Sharpe, E. J. Fox, A. W. Barnard, J. Finney, K. Watanabe, T. Taniguchi, M. Kastner, and D. Goldhaber-Gordon, Emergent ferromagnetism near three-quarters filling in twisted bilayer graphene, Science 365, 605 (2019).
- Nuckolls et al. [2020] K. P. Nuckolls, M. Oh, D. Wong, B. Lian, K. Watanabe, T. Taniguchi, B. A. Bernevig, and A. Yazdani, Strongly correlated chern insulators in magic-angle twisted bilayer graphene, Nature , 1 (2020).
- Wu et al. [2021] S. Wu, Z. Zhang, K. Watanabe, T. Taniguchi, and E. Y. Andrei, Chern insulators, van hove singularities and topological flat bands in magic-angle twisted bilayer graphene, Nature materials 20, 488 (2021).
- Choi et al. [2021] Y. Choi, H. Kim, Y. Peng, A. Thomson, C. Lewandowski, R. Polski, Y. Zhang, H. S. Arora, K. Watanabe, T. Taniguchi, et al., Correlation-driven topological phases in magic-angle twisted bilayer graphene, Nature 589, 536 (2021).
- Xie et al. [2021] Y. Xie, A. T. Pierce, J. M. Park, D. E. Parker, E. Khalaf, P. Ledwith, Y. Cao, S. H. Lee, S. Chen, P. R. Forrester, et al., Fractional chern insulators in magic-angle twisted bilayer graphene, Nature 600, 439 (2021).
- Park et al. [2023] H. Park, J. Cai, E. Anderson, Y. Zhang, J. Zhu, X. Liu, C. Wang, W. Holtzmann, C. Hu, Z. Liu, et al., Observation of fractionally quantized anomalous Hall effect, Nature 622, 74 (2023).
- Lu et al. [2023] Z. Lu, T. Han, Y. Yao, A. P. Reddy, J. Yang, J. Seo, K. Watanabe, T. Taniguchi, L. Fu, and L. Ju, Fractional quantum anomalous hall effect in a graphene moire superlattice, arXiv preprint arXiv:2309.17436 (2023).
- Saito et al. [2021] Y. Saito, J. Ge, L. Rademaker, K. Watanabe, T. Taniguchi, D. A. Abanin, and A. F. Young, Hofstadter subband ferromagnetism and symmetry-broken chern insulators in twisted bilayer graphene, Nature Physics 17, 478 (2021).
- Lin et al. [2022] J.-X. Lin, Y.-H. Zhang, E. Morissette, Z. Wang, S. Liu, D. Rhodes, K. Watanabe, T. Taniguchi, J. Hone, and J. Li, Spin-orbit–driven ferromagnetism at half moiré filling in magic-angle twisted bilayer graphene, Science 375, 437 (2022).
- Cao et al. [2021] Y. Cao, D. Rodan-Legrain, J. M. Park, N. F. Yuan, K. Watanabe, T. Taniguchi, R. M. Fernandes, L. Fu, and P. Jarillo-Herrero, Nematicity and competing orders in superconducting magic-angle graphene, science 372, 264 (2021).
- Cao et al. [2018a] Y. Cao, V. Fatemi, A. Demir, S. Fang, S. L. Tomarken, J. Y. Luo, J. D. Sanchez-Yamagishi, K. Watanabe, T. Taniguchi, E. Kaxiras, et al., Correlated insulator behaviour at half-filling in magic-angle graphene superlattices, Nature 556, 80 (2018a).
- Cao et al. [2018b] Y. Cao, V. Fatemi, S. Fang, K. Watanabe, T. Taniguchi, E. Kaxiras, and P. Jarillo-Herrero, Unconventional superconductivity in magic-angle graphene superlattices, Nature 556, 43 (2018b).
- Yankowitz et al. [2019] M. Yankowitz, S. Chen, H. Polshyn, Y. Zhang, K. Watanabe, T. Taniguchi, D. Graf, A. F. Young, and C. R. Dean, Tuning superconductivity in twisted bilayer graphene, Science 363, 1059 (2019).
- Khalaf et al. [2021] E. Khalaf, S. Chatterjee, N. Bultinck, M. P. Zaletel, and A. Vishwanath, Charged skyrmions and topological origin of superconductivity in magic-angle graphene, Science advances 7, eabf5299 (2021).
- Chatterjee et al. [2022] S. Chatterjee, M. Ippoliti, and M. P. Zaletel, Skyrmion superconductivity: DMRG evidence for a topological route to superconductivity, Phys. Rev. B 106, 035421 (2022).
- Törmä et al. [2022] P. Törmä, S. Peotta, and B. A. Bernevig, Superconductivity, superfluidity and quantum geometry in twisted multilayer systems, Nature Reviews Physics 4, 528 (2022).
- Ding et al. [2022] K. Ding, C. Fang, and G. Ma, Non-hermitian topology and exceptional-point geometries, Nature Reviews Physics 4, 745 (2022).
- Ashida et al. [2020] Y. Ashida, Z. Gong, and M. Ueda, Non-hermitian physics, Advances in Physics 69, 249 (2020).
- Bergholtz et al. [2021] E. J. Bergholtz, J. C. Budich, and F. K. Kunst, Exceptional topology of non-hermitian systems, Reviews of Modern Physics 93, 015005 (2021).
- Xu et al. [2017] Y. Xu, S.-T. Wang, and L.-M. Duan, Weyl exceptional rings in a three-dimensional dissipative cold atomic gas, Physical review letters 118, 045701 (2017).
- Cerjan et al. [2018] A. Cerjan, M. Xiao, L. Yuan, and S. Fan, Effects of non-hermitian perturbations on weyl hamiltonians with arbitrary topological charges, Physical Review B 97, 075128 (2018).
- Yao and Wang [2018] S. Yao and Z. Wang, Edge states and topological invariants of non-hermitian systems, Physical review letters 121, 086803 (2018).
- Kawabata et al. [2019] K. Kawabata, K. Shiozaki, M. Ueda, and M. Sato, Symmetry and topology in non-hermitian physics, Physical Review X 9, 041015 (2019).
- Gong et al. [2018] Z. Gong, Y. Ashida, K. Kawabata, K. Takasan, S. Higashikawa, and M. Ueda, Topological phases of non-hermitian systems, Physical Review X 8, 031079 (2018).
- Shao et al. [2024] K. Shao, H. Geng, E. Liu, J. L. Lado, W. Chen, and D. Xing, Non-hermitian moiré valley filter, Physical Review Letters 132, 156301 (2024).
- Zhang et al. [2019] Y.-R. Zhang, Z.-Z. Zhang, J.-Q. Yuan, M. Kang, and J. Chen, High-order exceptional points in non-hermitian moire lattices, Frontiers of Physics 14, 1 (2019).
- Ahn et al. [2019] J. Ahn, S. Park, and B.-J. Yang, Failure of nielsen-ninomiya theorem and fragile topology in two-dimensional systems with space-time inversion symmetry: Application to twisted bilayer graphene at magic angle, Phys. Rev. X 9, 021013 (2019).
- Nielsen and Ninomiya [1981] H. B. Nielsen and M. Ninomiya, No-go theorum for regularizing chiral fermions, Tech. Rep. (Science Research Council, 1981).
- Song et al. [2021] Z.-D. Song, B. Lian, N. Regnault, and B. A. Bernevig, Twisted bilayer graphene. ii. stable symmetry anomaly, Physical Review B 103, 205412 (2021).
- Carr et al. [2019] S. Carr, S. Fang, Z. Zhu, and E. Kaxiras, Exact continuum model for low-energy electronic states of twisted bilayer graphene, Physical Review Research 1, 013001 (2019).
- Tarnopolsky et al. [2019a] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Physical review letters 122, 106405 (2019a).
- Nam and Koshino [2017] N. N. T. Nam and M. Koshino, Lattice relaxation and energy band modulation in twisted bilayer graphene, Phys. Rev. B 96, 075311 (2017).
- Tarnopolsky et al. [2019b] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Phys. Rev. Lett. 122, 106405 (2019b).
- Shen et al. [2018] H. Shen, B. Zhen, and L. Fu, Topological band theory for non-hermitian hamiltonians, Physical review letters 120, 146402 (2018).
- Oudich et al. [2021] M. Oudich, G. Su, Y. Deng, W. Benalcazar, R. Huang, N. J. R. K. Gerard, M. Lu, P. Zhan, and Y. Jing, Photonic analog of bilayer graphene, Phys. Rev. B 103, 214311 (2021).
- Lou et al. [2021] B. Lou, N. Zhao, M. Minkov, C. Guo, M. Orenstein, and S. Fan, Theory for twisted bilayer photonic crystal slabs, Phys. Rev. Lett. 126, 136101 (2021).
- Sunku et al. [2018] S. Sunku, G. Ni, B.-Y. Jiang, H. Yoo, A. Sternbach, A. McLeod, T. Stauber, L. Xiong, T. Taniguchi, K. Watanabe, et al., Photonic crystals for nano-light in moiré graphene superlattices, Science 362, 1153 (2018).
- Deng et al. [2020] Y. Deng, M. Oudich, N. J. Gerard, J. Ji, M. Lu, and Y. Jing, Magic-angle bilayer phononic graphene, Physical Review B 102, 180304 (2020).
- Valagiannopoulos [2022] C. Valagiannopoulos, Electromagnetic analog to magic angles in twisted bilayers of two-dimensional media, Physical Review Applied 18, 044011 (2022).
- Oudich et al. [2022] M. Oudich, Y. Deng, and Y. Jing, Twisted pillared phononic crystal plates, Applied Physics Letters 120 (2022).
- Gardezi et al. [2021] S. M. Gardezi, H. Pirie, S. Carr, W. Dorrell, and J. E. Hoffman, Simulating twistronics in acoustic metamaterials, 2D Materials 8, 031002 (2021).
- Lou et al. [2022] B. Lou, B. Wang, J. A. Rodríguez, M. Cappelli, and S. Fan, Tunable guided resonance in twisted bilayer photonic crystal, Science Advances 8, eadd4339 (2022).
- Tang et al. [2023] H. Tang, B. Lou, F. Du, M. Zhang, X. Ni, W. Xu, R. Jin, S. Fan, and E. Mazur, Experimental probe of twist angle–dependent band structure of on-chip optical bilayer photonic crystal, Science Advances 9, eadh8498 (2023).
- Meng et al. [2023] Z. Meng, L. Wang, W. Han, F. Liu, K. Wen, C. Gao, P. Wang, C. Chin, and J. Zhang, Atomic bose–einstein condensate in twisted-bilayer optical lattices, Nature 615, 231 (2023).
- Cerjan et al. [2019] A. Cerjan, S. Huang, M. Wang, K. P. Chen, Y. Chong, and M. C. Rechtsman, Experimental realization of a weyl exceptional ring, Nature Photonics 13, 623 (2019).
- Song et al. [2023] W. Song, S. Wu, C. Chen, Y. Chen, C. Huang, L. Yuan, S. Zhu, and T. Li, Observation of weyl interface states in non-hermitian synthetic photonic systems, Physical Review Letters 130, 043803 (2023).
- Yu et al. [2024] X.-J. Yu, Z. Pan, L. Xu, and Z.-X. Li, Non-hermitian strongly interacting dirac fermions, Physical Review Letters 132, 116503 (2024).

### v2

@@
 In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss.
-Using a non-Hermitian generalization of Bistritzer-MacDonald(BM) model, we find Dirac cones centered at corner of the moiré Brillouin zone deformed in the presence of non-Hermiticity.
-This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both and corners of the Brillouion zone.
+Using a non-Hermitian generalization of Bistritzer-MacDonald model, we find Dirac cones centered at only () corner of the moiré Brillouin zone at () valley deformed in the presence of non-Hermiticity.
+This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both and corners of the Brillouin zone.
 The coincident of exceptional rings at point characterizes an “exceptional magic angle”, at which the system hosts flat bands with zero energy and finite lifetime.
@@
 Graphene owes extraordinary electronic and optical properties in two-dimensional (2D) systems [1, 2].
-These properties relate to its lattice structure, which is directly interconnected with its wave-function topology.
-In single-layer graphene, the presence of unusual linear band dispersion termed Dirac band dispersion at and points of the Brillouin zone is derived from its honeycomb lattice geometry.
-The Lorentz invariant Dirac cone at and points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4].
+These properties relate to its lattice structure, which is directly interconnected with its topological property.
+In single-layer graphene, the presence of unusual linear band dispersion termed “Dirac cone” at and points of the Brillouin zone is derived from its honeycomb lattice geometry.
+The two Dirac points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4].
 When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6].
@@
 The recent studies on twisted bilayer graphene (TBG) focus on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12].
-It is related to a variety of exotic phrases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29].
+It is related to a variety of exotic phases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29].
+On the other hand, twisted bilayer graphene has extraordinary topological properties.
+Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in two-dimensional periodic systems following the Nielsen-Ninomiya theorem [30, 31].
+This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges.
 The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [32, 33, 34].
@@
 Despite topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of twisted moiré system with magic angle has been overlooked [40, 41].
-In particular, twisted bilayer graphene has extraordinary topological properties.
-Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in 2D periodic systems following the Nielsen-Ninomiya theorem [40, 41].
-This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges.
-Actually, when the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [34].
-It is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
+Usually, non-Hermiticity is detrimental to the topology in single-layer graphene.
+When the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [36].
+In single-layer graphene, the annihilation of exceptional points from different valleys has been experimentally observed in a two-dimensional photonic system [42].
+Since the moire BZ valley of TBG has different topological properties from single-layer graphene, it is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
 In this paper, we uncover a new type of twisted band engineering in TBG, with non-Hermitian perturbation on one of the layers.
 As shown in Fig.
-1(a), balanced gain (in bright color) and loss (in light color) are put on the bottom layers (in red and blue, respectively).
+1(a), balanced gain and loss are put on the bottom layers (in red and blue, respectively).
 The superposing of two graphene layers with a twist angle between them creates moiré patterns.
@@
 1(b), the low-energy band structure consists of Dirac cones from the rotated layers located at the and corners.
-When the non-Hermitian perturbation is turned on the bottom layer, Dirac points morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig.
-1(c).
+When the non-Hermitian perturbation is turned on the bottom layer, Dirac points at () corner in the mBZ at () valley morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig.
+1(c) and (d) respectively.
 We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at points in an mBZ.
@@
 II The non-Hermitian bilayer model
-We consider balanced gain and loss on both layers of the bilayer system.
+We consider balanced gain and loss on the bottom layer of the bilayer system.
 We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene.
-The low-energy Hamiltonian for the mBZ is:
+The low-energy Hamiltonian for the mBZ at or valley is:
 where the top layer Hamiltonian is
@@
 Here, is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity.
-And corresponds to the rotated Pauli matrices .
+And corresponds to the rotated Pauli matrices with for and valley respectively.
 is the sublattice-dependent moiré potential that couples the two layers.
@@
 In the absence of non-Hermiticity, this is not a magic angle with a flat band.
-If the interlayer coupling effect is absent (), the dispersion in the moiré Brillouin zone is equivalent to the folded dispersion in the Brillouin zone of graphene.
-The Dirac point of the point in layer 1 is mapped to , while that of the point of layer 2 is mapped to .
+If the interlayer coupling effect is absent (), the dispersion in the mBZ is equivalent to the folded dispersion in the Brillouin zone of graphene.
+Since the physics in valley and valley are related, we only discuss the valley in the following.
+The Dirac point at the corner in the top layer is mapped to point in the mBZ, while that at the corner of the bottom layer is mapped to point in the mBZ.
 In the presence of non-Hermitian perturbation , the energy dispersion becomes complex.
@@
 The latter has deformed Dirac cones in both and corners in the Brillouin zone.
-This difference is due to the fact that and in non-Hermitian TBG are folded from different graphene layers.
+This difference is due to the fact that the Dirac cones at and point in non-Hermitian TBG are folded from different graphene layers.
 The non-Hermitian can only deform the Dirac cone on the bottom layer.
@@
 Thus, the Dirac cone is turned into a ring of exceptional points, this is similar to that in non-Hermitian semimetals [35].
-At a critical non-Hermitian strength, the exceptional point approach point, where the whole lowest band becomes flat, as shown in Fig.
+At a critical non-Hermitian strength , the exceptional point approach point, where the real part whole lowest band becomes flat, as shown in Fig.
 3.
-Different metric including the vanishing of Dirac velocity, maximum gap to neighboring bands, and minimum bandwidth gives the same results.
-The value of this critical non-Hermiticity strength depends on the twist angle .
+This characterizes an “exceptional magic” angle.
+According to the geometry relationship, is determined by the lattice size of mBZ and Fermi velocity .
 IV Exceptional cone
-To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. LABEL:eq:H at the first honeycomb shell. The Hamiltonian is written as
+To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. 1 at the first honeycomb shell. The Hamiltonian is written as
 where for are on the top layer and on the bottom layer.
@@
 It decreases with increasing interlayer coupling in AB region .
-To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases.
+To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases for and .
 When , increases with increasing interlayer coupling in AB region .
 Thus, the ratio becomes larger as increases.
-When, increases with increasing interlayer coupling in AA region .
-The enon-Hermitian effect is more complicated.
+When , increases with increasing interlayer coupling in AA region and the non-Hermitian effect is more complicated.
+We can expect that the exceptional magic angle is not only determined by but also .
 V Berry Curvature and Chern number
@@
 On the boundary of exceptional rings, the wave functions vary rapidly and will give a large Berry curvature.
-Upon summing the Berry curvature on the mmBZ, we can obtain the total Berry charge or Chern number
+Upon summing the Berry curvature on the mBZ, we can obtain the total Berry charge or Chern number
 The non-Hermitian perturbation transforms a Dirac point into an exceptional ring.
 In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners.
-We find that the total Berry charge in mBZ is 1 throughout the merging of exceptional rings with increasing (e.g., for ).
-At much larger , the Berry charge can be a quantized value larger than 1 due to approaching band crossings with the remote bands.
-This supports our expectation that the topological charge in the system is preserved in the exceptional contour.
-And most importantly, this characterizes the topological features of non-Hermitian TBG.
-This is opposite to the merging of exceptional contours with opposite topological charge, in which the topological charge is dissipated and gives a single, uncharged exceptional contour.
+For small , we can find that the total Berry charge in one mBZ is 1.
+This can be easy to understand from two facts.
+One is that each of the three Dirac points corresponds to Berry curvature monopoles of charge 1 in the absence of non-Hermiticity.
+The other is that the topological charge is preserved on the exceptional contour that forms from the original Dirac point.
+Thus, we can expect that each of the three exceptional contours contributes one-thirds of the Berry charge to one mBZ.
+Most interestingly, even when exceeds , the total Berry charge remains one.
+This is opposite to the merging of exceptional contours with opposite topological charge [36], in which the topological charge is dissipated and gives a single, uncharged exceptional contour.
+This indicates a breakdown of Nielsen-Ninomiya theorem and distinguishes the valley topological features of non-Hermitian TBG from a single-layer graphene.
+At much larger (e.g., for ), the Berry charge can be a quantized value larger than 1 due to the band crossings with the remote bands.
 VI Discussion and conclusion
 In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers.
-We found that exceptional rings centered at corner in the moiré Brillouin zone in the () valley.
+We found that exceptional rings centered at () corner in the moiré Brillouin zone in the () valley.
 This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene.
@@
 Note added-.
-Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed.
+Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed [61].
 Acknowledgements.
 This work is supported by the National Natural Science Foundation of China (Grant No. 12104099 and No. 12274095) and Guangzhou Science and Technology Program (Grant No. 2024A04J0272)References
-- Neto et al. [2009] A. C. Neto, F. Guinea, N. M. Peres, K. S. Novoselov, and A. K. Geim, The electronic properties of graphene, Reviews of modern physics 81, 109 (2009).
+- CastroNeto et al. [2009] A. H. CastroNeto, F. Guinea, N. M. R. Peres, K. S. Novoselov, and A. K. Geim, The electronic properties of graphene, Reviews of modern physics 81, 109 (2009).
 - Gonçalves and Peres [2016] P. A. D. Gonçalves and N. M. Peres, An introduction to graphene plasmonics (World Scientific, 2016).
@@
 - Charlier et al. [1992] J.-C. Charlier, J.-P. Michenaud, and P. Lambin, Tight-binding density of electronic states of pregraphitic carbon, Phys. Rev. B 46, 4540 (1992).
-- McCann [2006] E. McCann, Asymmetry gap in the electronic band structure of bilayer graphene, Phys. Rev. B 74, 161403 (2006).
+- McCann [2006] E. McCann, Asymmetry gap in the electronic band structure of bilayer graphene, Phys. Rev. B 74, 161403(R) (2006).
 - Dos Santos et al. [2007] J. L. Dos Santos, N. Peres, and A. C. Neto, Graphene bilayer with a twist: electronic structure, Physical review letters 99, 256802 (2007).
 - Bistritzer and MacDonald [2011a] R. Bistritzer and A. H. MacDonald, Moiré bands in twisted double-layer graphene, Proceedings of the National Academy of Sciences 108, 12233 (2011a).
-- Bistritzer and MacDonald [2011b] R. Bistritzer and A. MacDonald, Moiré butterflies in twisted bilayer graphene, Physical Review B 84, 035440 (2011b).
+- Bistritzer and MacDonald [2011b] R. Bistritzer and A. H. MacDonald, Moiré butterflies in twisted bilayer graphene, Physical Review B 84, 035440 (2011b).
 - Dos Santos et al. [2012] J. L. Dos Santos, N. Peres, and A. C. Neto, Continuum model of the twisted graphene bilayer, Physical Review B 86, 155449 (2012).
@@
 Kong, R.
+R.
 Nair, K.
@@
 - Törmä et al. [2022] P. Törmä, S. Peotta, and B. A. Bernevig, Superconductivity, superfluidity and quantum geometry in twisted multilayer systems, Nature Reviews Physics 4, 528 (2022).
+- Ahn et al.
+[2019] J.
+Ahn, S.
+Park, and B.-J.
+Yang, Failure of nielsen-ninomiya theorem and fragile topology in two-dimensional systems with space-time inversion symmetry: Application to twisted bilayer graphene at magic angle, Phys.
+Rev.
+X 9, 021013 (2019).
+- Nielsen and Ninomiya [1981] H. B. Nielsen and M. Ninomiya, No-go theorum for regularizing chiral fermions, Tech. Rep. (Science Research Council, 1981).
 - Ding et al. [2022] K. Ding, C. Fang, and G. Ma, Non-hermitian topology and exceptional-point geometries, Nature Reviews Physics 4, 745 (2022).
@@
 - Gong et al. [2018] Z. Gong, Y. Ashida, K. Kawabata, K. Takasan, S. Higashikawa, and M. Ueda, Topological phases of non-hermitian systems, Physical Review X 8, 031079 (2018).
-- Shao et al. [2024] K. Shao, H. Geng, E. Liu, J. L. Lado, W. Chen, and D. Xing, Non-hermitian moiré valley filter, Physical Review Letters 132, 156301 (2024).
+- Shao et al. [2024] K. Shao, H. Geng, E. Liu, J. L. Lado, W. Chen, and D. Y. Xing, Non-hermitian moiré valley filter, Physical Review Letters 132, 156301 (2024).
 - Zhang et al. [2019] Y.-R. Zhang, Z.-Z. Zhang, J.-Q. Yuan, M. Kang, and J. Chen, High-order exceptional points in non-hermitian moire lattices, Frontiers of Physics 14, 1 (2019).
-- Ahn et al.
-[2019] J.
-Ahn, S.
-Park, and B.-J.
-Yang, Failure of nielsen-ninomiya theorem and fragile topology in two-dimensional systems with space-time inversion symmetry: Application to twisted bilayer graphene at magic angle, Phys.
-Rev.
-X 9, 021013 (2019).
-- Nielsen and Ninomiya [1981] H. B. Nielsen and M. Ninomiya, No-go theorum for regularizing chiral fermions, Tech. Rep. (Science Research Council, 1981).
+- Król et al.
+[2022] M.
+Król, I.
+Septembre, P.
+Oliwa, M.
+Kkedziora, K.
+Łempicka-Mirek, M.
+Muszyński, R.
+Mazur, P.
+Morawiak, W.
+Piecek, P.
+Kula, et al., Annihilation of exceptional points from different dirac valleys in a 2d photonic system, Nature Communications 13, 5340 (2022).
 - Song et al. [2021] Z.-D. Song, B. Lian, N. Regnault, and B. A. Bernevig, Twisted bilayer graphene. ii. stable symmetry anomaly, Physical Review B 103, 205412 (2021).
+- Tarnopolsky et al. [2019a] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Phys. Rev. Lett. 122, 106405 (2019a).
 - Carr et al. [2019] S. Carr, S. Fang, Z. Zhu, and E. Kaxiras, Exact continuum model for low-energy electronic states of twisted bilayer graphene, Physical Review Research 1, 013001 (2019).
-- Tarnopolsky et al. [2019a] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Physical review letters 122, 106405 (2019a).
-- Nam and Koshino [2017] N. N. T. Nam and M. Koshino, Lattice relaxation and energy band modulation in twisted bilayer graphene, Phys. Rev. B 96, 075311 (2017).
-- Tarnopolsky et al. [2019b] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Phys. Rev. Lett. 122, 106405 (2019b).
+- Tarnopolsky et al. [2019b] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Physical review letters 122, 106405 (2019b).
 - Shen et al. [2018] H. Shen, B. Zhen, and L. Fu, Topological band theory for non-hermitian hamiltonians, Physical review letters 120, 146402 (2018).
+- Meng et al. [2023] Z. Meng, L. Wang, W. Han, F. Liu, K. Wen, C. Gao, P. Wang, C. Chin, and J. Zhang, Atomic bose–einstein condensate in twisted-bilayer optical lattices, Nature 615, 231 (2023).
 - Oudich et al. [2021] M. Oudich, G. Su, Y. Deng, W. Benalcazar, R. Huang, N. J. R. K. Gerard, M. Lu, P. Zhan, and Y. Jing, Photonic analog of bilayer graphene, Phys. Rev. B 103, 214311 (2021).
@@
 Mazur, Experimental probe of twist angle–dependent band structure of on-chip optical bilayer photonic crystal, Science Advances 9, eadh8498 (2023).
-- Meng et al. [2023] Z. Meng, L. Wang, W. Han, F. Liu, K. Wen, C. Gao, P. Wang, C. Chin, and J. Zhang, Atomic bose–einstein condensate in twisted-bilayer optical lattices, Nature 615, 231 (2023).
 - Cerjan et al. [2019] A. Cerjan, S. Huang, M. Wang, K. P. Chen, Y. Chong, and M. C. Rechtsman, Experimental realization of a weyl exceptional ring, Nature Photonics 13, 623 (2019).
@@
 - Yu et al. [2024] X.-J. Yu, Z. Pan, L. Xu, and Z.-X. Li, Non-hermitian strongly interacting dirac fermions, Physical Review Letters 132, 116503 (2024).
+- Esparza and Juricic [2024] J. P. Esparza and V. Juricic, Exceptional magic angles in non-hermitian twisted bilayer graphene, arXiv preprint arXiv:2408.08804 (2024).